# CayleyPy RL: Pathfinding and Reinforcement Learning on Cayley Graphs

### v1

@@ 4. Mathematical contributions
In the section we prove lower and upper bounds for the diameter and for particular elements, describe decomposition of the conjecturally longest element of the desired length $n(n-1)/2$, describe algorithms to decompose elements in LRX generators and formulate several conjectures on the growth, random walks and spectrum of the LRX Cayley graph.
4.1. LRX Cayley graph
LRX generators are: L - left cyclic shift, R - right cyclic shift, X - eXchange (flip) of positions 0,1. It can be considered as a relative of the TopSpin and Hungarian Ring puzzles (figure 3). Despite its simplicicity its structure is not fully understood. In particular according to OEIS-A186783 the diameter conjecture $n(n-1)/2$ remains open.
Figure 1 is a visualization of the LRX Cayley graph for $S_{5}$ which shows its complicated structure. More visualizations by the minimal distortion method can be found at notebook, and by originally modified force directed layout here.
Figure 3. LRX and related mechanical puzzles
@@ 4.2. Long and the longest element
Figure 4. Permutations arising from dihedral symmetries - are provebly long elements. The longest is conjecturally one of them. Surprisingly it is not the full inversion.
We created efficient implementation of the brute force breadth first search algorithm notebook using bit operations over int64 and were able to traverse all the permutations for $n\leq 14$, outperforming OEIS data by one step. From that computation we observed that the longest element is unique and have a specific form. It can be described as a symmetry of a regular polygon (figure 4), which is orthogonal to 0-1 edge. Or explicitly as product of transpositions $(0,1)(n-1,2)(n-2,3)(n-3,3)...$. We conjecture that this element is unique longest element and it has length $n(n-1)/2$. Surprisingly not the full inversion ($i\to n-i-1$) is the longest.
As we will prove below actually all the permutations which correspond to symmetries of a regular polygon are also long elements.
@@
The element $(0,1)(n-1,2)(n-2,3)(n-3,3)...$ (conjecturally the longest) has the following explicit decomposition into product of LRX generators:
$\prod_{i=1}^{[(n-1)/2]}(XL^{(-1)^{(i-1)}})^{i}\prod_{i=([n/2]-1)}^{1}(L^{(-1)^{i+[(n-1)/2]-[n/2]}}X)^{i}R^{[n/2]}$ (4.1)
@@ 4.3. Lower bound on the diameter
Below, we provide a lower bound on the diameter using a combinatorial argument. The key idea is that cyclic shifts preserve the cyclic order of any given triple of numbers. Therefore, for each triple where the order is disrupted, t $X$ is required in the decomposition. The symmetries of a regular polygon, as described above, correspond precisely to the permutations that maximize changes in cyclic order. (The code to generate these long elements: notebook.)
@@
Diameter of LRX graph is larger or equal $n(n-1)/2-n/2-1$
At first, we reformulate the setup as follows.
The numbers $1,\ldots,n$ are written in the vertices of a regular $n$-gon $A_{1}A_{2}\ldots A_{n}$, each number being written once. At each move, one can either rotate the arrangement by $2\pi/n$ counterclockwise or clockwise, or switch the numbers at $A_{1}$ and $A_{2}$.
For two arrangements $\pi_{1},\pi_{2}$ denote by $d(\pi_{1},\pi_{2})$ the minimal number of moves which suffice to get $\pi_{2}$ from $\pi_{1}$.
@@
Proof. Assume that $\pi_{2}$ is obtained from $\pi_{1}$ by several moves. Call two numbers $a,b\in\{1,\ldots,n\}$ friends, if they were switched odd number of times. Since for every three numbers $a,b,c$ the orientation of a triangle formed by these numbers changed, there is odd number of pairs of friends between $a,b,c$. Let $A$ denote the set of all non-friends of element 1, $B$ the complement of $A$, i. e. $B$ consists of 1 and its friends. Our condition yields that all elements of $A$ are mutual friends, and so are all elements of $B$, but there is no friendship between $A$ and $B$. Thus, the total number of pairs of friends is $n(n-1)/2-|A|\cdot|B|\geqslant n(n-1)/2-n^{2}/4$. Thus there were at least as many switches. Between every two switches there was a rotation (otherwise we do something useless). Totally, the number of operations is not less than $n(n-1)-n^{2}/2-1=n^{2}/2-n-1$.
@@ 4.4. Upper bound on the diameter and the algorithms
We develop two algorithms which can decompose any permutation in the product of LRX generators. For the first we prove complexity bound $n(n-1)/2+3n$, thus bounding diameter from above. The code and tests for the first algorithm can be found in notebook.
@@
3) The element $a_{1}$ is placed in the $a_{n}$-th position, and we perform a sequence of elementary transpositions to move $a_{1}$ from the $a_{n}$-th position to the $a_{1}$-th position. During each elementary transposition, we either increase or decrease $x$ by 1. The sign of the change depends on the direction of the elementary transposition. Essentially, $x$ represents the current position of the transposition, but if our transposition is $(1,N)$, the change in $x$ is also ±$\pm 1$, even though the position changes from 1 to $N$ or from $N$ to 1.
@@
Complexity estimation. First, when decomposing a cycle into transpositions, we do not shift anything by more than 1, except for elements that have already been affected by the cycle. Second, when moving an element from position $i$ (or $i\pm 1$) to position $\sigma(i)$, we perform a sequence of at most $\text{dist}(i,\sigma(i))+1$ transpositions. Third, the transition from $i\to\sigma(i)$ to $\sigma(i)\to\sigma(\sigma(i))$ takes at most one rotation. Thus, we perform at most
@@
We also develop a second algorithm: notebook. Currently it is our top performing algorithm from the practical tests, we expect its complexity is bounded by $n(n-1)/2+n/2$, however it is not yet proved.
@@
Based on explicit computations for the growth for $n\leq 14$ and its analysis (notebooks 1 2 ) we come to the following conjecture, which can be thought as an analogue of the central limit theorems. It is in the vein of works by P.Diaconis et.al.  [Diaconis77, Diaconis88, Diaconis17] where for Coxeter generators growth and some other statistics were shown to follow Gaussian normal law in a limit. New point of our conjecture - appearance of asymmetric Gumbel distribution. Such asymmetry is quite natural - for the randomly chosen generators growth will be highly asymmetric with exponential growth and sharp abrupt (Rubik’s cube growth is a prototypical example), while for Coxeter generators which are nearly commutative - growth is symmetric Gaussian close to commutative case. Slight asymmetry of growth for LRX generators indicates its intermediate nature between ”exponential growth”/”commutative-like Gaussian growth” cases.
The growth - Gumbel. The growth distribution of the LRX graph follows the Gumbel distribution for large $n$.
@@
We have computed all possible shortest paths from the conjecturally longest element to the identity by methods of the dynamical programming (notebook ). And after that analyzed averaged trajectories (notebook) - which are analogs of random sorting networks [Virág]. The figure 6 below represents the results. It seems the pattern is different from the ”sine curves” discovered in [Virág].
@@
P.Diacons and L.Soloff-Coste studied random walks on LRX Cayley [Diaconis93]. They estimated that mixing time is between $n^{2}$ and $n^{3}log(n)$. We performed numerical experiments to make more precise conjecture (notebooks 1,2) unfortunately the achieved range of $n$ was not enough to draw conclusions, however it seems that mixing time is more close to $n^{3}log(n)$, moreover even for non-backtracking random walks, which is a bit surprising since they typically mix faster.
Finally we considered the Schreier coset graph for the action of $S_{2n}$ on binary strings with $n$ zeros and $n$ ones. (Analogue of $Gr(n,2n)$ over the ”field with one element”). Efficient code which stores such strings in int64 allowed us to completely traverse the graph up to $n=42$ and compute its growth. The analysis of the data presented in the notebook. Based on that it is tempting to conjecture that the diameters of such graphs are between $n^{2}/6$ and $n^{2}/5$, while growth again can be approximated by the Guimbel distribution.
@@ 5. Code availability
The link above leads to one of the three public Kaggle challenges which we created to stimulate the research and interplay between artificial intelligence and mathematical communities, the other two: LRX discover math and God’s algorithm, LRX-binary: in search of God’s number. Kaggle infrastructure provides a convenient way to benchmark and compare different algorithms - just making the submissions to these challenges. The code, the data and discussions are stored at the same place. Code can be executed for free on Kaggle cloud servers. The first challenge asks to decompose the conjecturally longest elements and to find their shortest decompositions - participants already achieved the length $n(n-1)/2$ an so if believe in the conjecture the optimum is found, if we do not believe - not. The second challenges asks to decompose random permutations and aims to find optimal algorithms (at the moment of writing optimum is not achieved), the third one asks the same but for the binary strings acted by LRX generators.
@@ Acknowledgements
A.C. is deeply grateful to M. Douglas for his interest in this work, engaging discussions, and invitation to present preliminary results at the Harvard CMSA program on ”Mathematics and Machine Learning” in Fall 2024, to M. Gromov, S. Nechaev, and V. Rubtsov for their invitation to make a talk (video) at ”Representations, Probability, and Beyond: A Journey into Anatoly Vershik World” at IHES, as well as for stimulating discussions. A.C. is grateful to J. Mitchel for involving into the Kaggle Santa 2023 challenge, from which this project originated and to M.Kontsevich, Y.Soibelman, S.Gukov, A. Hayat, T. Smirnova-Nagnibeda, D.Osipov, V. Kleptsyn, G.Olshanskii, A.Ershler, J. Ellenberg, G. Williamson, A. Sutherland, Y. Fregier, P.A. Melies, I. Vlassopoulos, F.Khafizov, A.Zinovyev, H.Isambert for the discussions, interest and comments, to his wife A.Chervova and daugther K.Chervova for support, understanding and help with computational experiments.

### v2

@@ 4. Mathematical contributions
In this section, we prove lower and upper bounds for the diameter and for particular elements, describe the decomposition of the conjecturally longest element of the desired length n(n1)/2n(n-1)/2, describe algorithms to decompose in LRX generators, and formulate several conjectures on the growth, random walks and spectrum of the LRX Cayley graph.
@@ 4.1. Long and the longest element
Figure 4. Permutations arising from dihedral symmetries are provably long elements. The longest is conjecturally one of them.
We created an efficient implementation of the brute force breadth first search algorithm (see this notebook) using bit operations over int64 and were able to traverse all the permutations for $n\leq 15$, outperforming OEIS data by two steps (for $n=15$ computation took several days on Hadoop cluster). From that computation we observed that the longest element is unique and has a specific form (1). It can be described as a symmetry of a regular polygon (Fig. 4), which is orthogonal to 0–1 edge. We conjecture that this element is the only longest element with length n(n1)/2n(n-1)/2. Surprisingly, the full inversion $(n-1,n-2,\ldots,2,1,0)$ is not the longest. As we will prove below, actually all the permutations which correspond to symmetries of a regular polygon are also long elements.
@@
Let
$n>2,\quad m=\Big\lfloor\frac{n}{2}\Big\rfloor,\quad\delta_{n}=[n=2m]=\begin{cases}1,n\text{ \rm{is} even},\\ 0,n\text{ \rm{is} odd}.\end{cases}$
The element $\ell_{n}=(1\ 0\ n-1\ n-2\ldots 3\ 2)$ has the following explicit decomposition into product of LRX generators:
$\ell_{n}=\prod_{i=1}^{m-\delta_{n}}\big(XL^{(-1)^{i-1}}\big)^{i}\circ\prod_{i=m-1}^{1}\big(L^{(-1)^{i-\delta_{n}}}X\big)^{i}\circ R^{m}.$ (5)
@@ 4.2. Lower bound on the diameter
Below, we provide a lower bound on the diameter using a combinatorial argument. The key idea is that cyclic shifts preserve the cyclic order of any given triple of numbers. Therefore, for each triple where the order is disrupted, the move $X$ is required in the decomposition. The symmetries of a regular polygon, as described above, correspond precisely to the permutations that maximize changes in cyclic order. (The code to generate these long elements: notebook.)
@@
The diameter of the LRX Cayley graph of $S_{n}$ is larger than or equal to n(n1)/2n/21n(n-1)/2-n/2-1.
First, we reformulate the setup as follows.
The numbers $1,\ldots,n$ are written in the vertices of a regular $n$-gon $A_{1}A_{2}\ldots A_{n}$, each number being written once. At each move, one can either rotate the arrangement by $2\pi/n$ counterclockwise or clockwise, or swap the numbers at $A_{1}$ and $A_{2}$. For two arrangements $\pi_{1},\pi_{2}$ denote by $d(\pi_{1},\pi_{2})$ the minimal number of moves which suffice to get $\pi_{2}$ from $\pi_{1}$.
@@
Proof. Assume that $\pi_{2}$ is obtained from $\pi_{1}$ by several moves. Call two numbers $a,b\in\{1,\ldots,n\}$ friends, if they were swapped odd number of times. Since for every three numbers $a,b,c$ the orientation of a triangle formed by these numbers changed, there is odd number of pairs of friends between $a,b,c$. Let $A$ denote the set of all non-friends of element 1, $B$ the complement of $A$, i. e. $B$ consists of 1 and its friends. Our condition yields that all elements of $A$ are mutual friends, and so are all elements of $B$, but there is no friendship between $A$ and $B$. Thus, the total number of pairs of friends is $n(n-1)/2-|A|\cdot|B|\geqslant n(n-1)/2-n^{2}/4$. Hence, there were at least as many swaps. Between any two swaps there should be a rotation (otherwise, we are doing something useless). Totally, the number of operations is not less than $n(n-1)-n^{2}/2-1=n^{2}/2-n-1$.
@@ 4.3. Upper bound on the diameter and the algorithms
We develop two algorithms which can decompose any permutation in the product of LRX generators. We prove complexity bound n(n1)/2+3nn(n-1)/2+3n, thus bounding diameter from above. The code and tests for the first algorithm can be found in notebook.
@@
3) The element $a_{1}$ is placed in the $a_{n}$-th position, and we perform a sequence of elementary transpositions to move $a_{1}$ from the $a_{n}$-th position to the $a_{1}$-th position. During each elementary transposition, we either increase or decrease $x$ by 1. The sign of the change depends on the direction of the elementary transposition. Essentially, $x$ represents the current position of the transposition, but if our transposition is $(1,N)$, the change in $x$ is also $\pm 1$, even though the position changes from 1 to $N$ or from $N$ to 1.
@@
Complexity estimation. First, when decomposing a cycle into transpositions, we do not shift anything by more than 1, except for elements already affected by the cycle. Second, when moving an element from position $i$ (or $i\pm 1$) to position $\sigma(i)$, we perform a sequence of at most $\text{dist}(i,\sigma(i))+1$ transpositions. Third, the transition from $i\to\sigma(i)$ to $\sigma(i)\to\sigma(\sigma(i))$ takes at most one rotation. Thus, we perform at most
@@
We also develop a second algorithm: notebook. Currently, it is our top performing algorithm from the practical tests, we expect its complexity is bounded by n(n1)/2+n/2n(n-1)/2+n/2, however, it is not yet proved.
@@
Based on explicit computations for the growth for $n\leq 15$ and its analysis (notebooks 1 2 ) we come to the following conjecture, which can be thought as an analogue of the central limit theorems. It is in the vein of works by P.Diaconis et.al.  [Diaconis77, Diaconis88, Diaconis17] where for Coxeter generators growth and some other statistics were shown to follow Gaussian normal law in a limit. A new point of our conjecture is the appearance of asymmetric Gumbel distribution. Such asymmetry is quite natural since for the randomly chosen generators growth will be highly asymmetric with exponential growth and sharp abrupt (Rubik’s cube growth is a prototypical example), while for nearly commutative Coxeter generators the growth is symmetric Gaussian. A slight asymmetry of growth for LRX generators indicates its intermediate nature between ”exponential growth” and ”commutative-like Gaussian growth” cases.
The growth distribution of the LRX graph seems to follow the Gumbel distribution for large $n$. The detailed analysis is provided in supplementary section 6.1.
@@
We have computed all possible shortest paths from the conjecturally longest element to the identity by methods of the dynamical programming (notebook . And after that analyzed averaged trajectories (notebook) which are analogs of random sorting networks [Virág]. The figure 6 below represents the results. It seems the pattern is different from the ”sine curves” discovered in [Virág].
@@ 5. Code availability
The link above leads to one of the three public Kaggle challenges which we created to stimulate the research and interplay between artificial intelligence and mathematical communities, the other two: LRX discover math and God’s algorithm, LRX-binary: in search of God’s number. Kaggle infrastructure provides a convenient way to benchmark and compare different algorithms by just making the submissions to these challenges. The code, the data and discussions are stored at the same place. The code can be executed for free on Kaggle cloud servers. The first challenge is to decompose the conjecturally longest elements and to find their shortest decompositions. The participants have already achieved the length n(n1)/2n(n-1)/2; thus, if we believe the conjecture is true, the optimum is found. The second challenge is to decompose random permutations and aims to find optimal algorithms (at the moment of writing optimum is not achieved). The third one asks the same but for the binary strings acted by LRX generators.
@@ Acknowledgements
A.C. is deeply grateful to M. Douglas for his interest in this work, engaging discussions, and invitation to present preliminary results at the Harvard CMSA program on ”Mathematics and Machine Learning” in Fall 2024; to M. Gromov, S. Nechaev, and V. Rubtsov for their invitation to give a talk (video) at ”Representations, Probability, and Beyond: A Journey into Anatoly Vershik World” at IHES, as well as for stimulating discussions. A.C. is grateful to J. Mitchel for involving into the Kaggle Santa 2023 challenge, from which this project originated, and to M.Kontsevich, Y.Soibelman, S.Gukov, A. Hayat, T. Smirnova-Nagnibeda, D.Osipov, V. Kleptsyn, G.Olshanskii, A.Ershler, J. Ellenberg, G. Williamson, A. Sutherland, Y. Fregier, P.A. Melies, I. Vlassopoulos, F.Khafizov, A.Zinovyev, H.Isambert, T. Rokicki for the discussions, interest and comments, to his wife A.Chervova and daugther K.Chervova for support, understanding and help with computational experiments.